OpenStudy (anonymous):

I have the answer can someone just see if I am correct? Rewrite the following function in the form f(x)=kx^p. f(x)=(4x^3)^2 k= p=

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

k=16 and p=5

OpenStudy (jtvatsim):

check your value for p

OpenStudy (anonymous):

is p=6

OpenStudy (anonymous):

k=16 p=6

OpenStudy (jtvatsim):

that is correct!

OpenStudy (anonymous):

Ok thank you

OpenStudy (jtvatsim):

x^3 + x^2 = x^5 but (x^3)^2 = x^6

OpenStudy (anonymous):

you should use superposition, principle to compare the alike

OpenStudy (anonymous):

oh ok